FORMATION OF EXCITED H_{3}O STATES BY CHARGE TRANSFER FROM Cs

\underline{STANISLAV A. DOLGIKH}, ANNA I. KRYLOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90007.

Charge transfer (CT) probability between Cs and H_{3}O^{+} was studied. In the experiment a Rydberg type H_{3}O molecule is formed in electronically excited state after CT from Cs, following by the dissociation to H_{2}O+H. However it is not clear which excited state is a precursor of the dissociation. Electronic structure of H_{3}O molecule was investigated by ab initio methods. Couplings between initial and target states were calculated, and charge transfer probabilities were estimated. Calculations show that 4s Rydberg state of H_{3}O molecule is the most likely state to be populated in the experiment.